OpenStudy (anonymous):

A rectangular prism is truncated removing 20% of its height. The original height of the prism was 100 inches and it has a length of 5 inches and a width of 10 inches. What is the volume of the truncated prism? 4,000 in3 2,000 in3 3,000 in3 5,000 in3

OpenStudy (anonymous):

V = L * W * H V = 5 * 10 * 100 V= 5000 in^3 original prism. Can you solve for the truncated one now? 100 * .8 = 20% of height....

OpenStudy (anonymous):

Rectangular prism..not a triangular prism. :-)